For this problem, carry at least four digits after the decimal in your calculations. Answers may vary slightly due to rounding.

A random sample of 5940 physicians in Colorado showed that 3170 provided at least some charity care (i.e., treated poor people at no cost).

(a) Let p represent the proportion of all Colorado physicians who provide some charity care. Find a point estimate for p. (Round your answer to four decimal places.)


(b) Find a 99% confidence interval for p. (Round your answers to three decimal places.)

lower limit    
upper limit    
Answer #1

Solution - given that, X = 3170 n = 5940 P = x = 3170 = 0.5337 5940 a point estimate P = 0.5337 - P = -0.5331 - 0.4663 b) is